2^z=8
thank you (:

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)
Answer by Alan3354(69443)

2^z=8
-----------
If you recognize that 8 = 2^3:
2^z = 2^3
x = 3
--------
If you don't:
z*log(2) = log(8)
z = log(8)/log(2)
z = 3